LASER SPECTROSCOPIC STUDIES OF CARBON MOLECULES

\underline{J. DUDEK}, M. C. MCCARTHY and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and Division of Engineering Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138.

Recent experiments to measure the electronic spectra of reactive carbon molecules of astrophysical interest will be presented. The molecules are generated using an electrical discharge through a precursor gas in the throat of a supersonic nozzle. Optical transitions are detected by one of three complementary methods: (1) Cavity Ring-down Spectroscopy (CRDS), (2) Laser Induced Fluorescence (LIF), and most recently (3) Resonant Two Color Two Photon Ionization (R2C2PI) with Time of Flight Mass Spectrometry (TOF-MS). Several small carbon chains and silicon-carbon rings have now been detected by these methods, allowing systematic comparisons of detection sensitivity and these results will be summarized. Finally, further results concerning the carrier of a strong laboratory band at λ4430 will be presented.